# StutterZero and StutterFormer: End-to-End Speech Conversion for Stuttering Transcription and Correction

### v1

@@ 1 Introduction
Stuttering is a prevalent speech disorder that affects more than 70 million individuals worldwide [1]. It manifests as interruptions in speech flow, including unintentional repetitions, prolongations, and pauses. Table 1 presents the five main phonological classifications of stuttering and examples of their symptoms [2] [3]. Such interruptions can significantly hinder effective communication, frequently causing social anxiety, depression, and a diminished quality of life [4]. In children, stuttering may lead to bullying and social isolation, exacerbating the emotional and psychological difficulties they encounter [5].
@@
Fluent speech plays a critical role in daily communication, both in interpersonal situations and when engaging with intelligent voice-based technologies. Through a quality-of-life survey, people who stutter (PWS) showed a statistically significant decrease in emotional health and social function metrics [6]. Stuttering can also cause feelings of shame, fear, anxiety, and guilt [7]. These challenges are exacerbated by the growing dependence on intelligent voice assistants, such as Google Home, Amazon Echo, and Apple Siri, which are designed to process fluent speech. As a result, people who stutter frequently encounter recognition errors, limited functionality, and exclusion when interacting with voice-controlled technologies.
For more severe forms of stuttering, automatic speech recognition (ASR) models, which transcribe speech, may insert unintended words or even truncate the speech due to a blocking stutter. Lea et al. [8] evaluated the speech of individuals with PWS using the Apple Speech framework, a production-level automatic speech recognition (ASR) system designed for fluent speakers. They reported a Word Error Rate (WER) of 19.8%, indicating that 19.8% of words in the ASR-generated transcript did not match the reference ground truth. They also reported a truncation rate of 23.8%, where 23.8% of utterances from these PWS were prematurely cut off [8]. Addressing this challenge, this study develops deep learning models that convert stuttered speech into fluent audio, going beyond transcription to reconstruct corrected acoustic signals. By generating fluent speech that preserves semantic content and improves intelligibility, these models aim to enhance communication for PWS and increase accessibility in human-machine interactions.
@@ 1.0.1 Conventional Digital Signal Processing Approaches
DSP approaches utilize audio feature extraction methods to obtain condensed numerical features from complex audio signals. Then, a ruleset or a set of predetermined filters is applied to the features to determine which timeframes contain a stutter. Finally, these timeframes are cut out of the audio, removing the stutter. Some frequently utilized feature sets include Mel-Frequency Cepstral Coefficients (MFCCs), Linear Predictive Coding (LPC), and Linear Predictive Cepstral Coefficients (LPCCs) [9]. For instance, MFCC features are generated by first applying the Fast Fourier Transform (FFT) to convert an audio sample from the amplitude domain to the frequency domain, generating the power spectrum. After that, the Mel filter bank is used to map the power spectrum to Mel frequencies, employing a set of nonlinear triangular filters. This aligns the intensity of frequencies to match the nonlinearity of human auditory perception. Finally, a Discrete Cosine Transform (DCT) is used to generate the cepstral coefficients through the decorrelation of features. This pipeline is applied to every window of audio, usually 20–50 ms long [9] [10].
K. N. et al. introduced a rule-based approach for stutter detection and removal by computing a “correlation factor” between adjacent audio windows using either MFCC or LPC features [10]. A high correlation value, empirically determined to be 0.92, was used to identify repeated or prolonged segments, prompting the deletion of redundant frames. The unusually high correlation factor was used to detect repeated audio patterns, such as recurring phonemes or extended silences. However, this approach was evaluated exclusively on repetition and prolongation stutters, without accounting for other common disfluencies such as blocks or interjections. Additionally, the experimental scope of the study was confined to only 60 repetition and 70 prolongation events obtained from a limited selection of audio files. As a result, its generalizability across diverse speakers, accents, or spontaneous conversational settings remains uncertain [10]. Table 2 summarizes the performance, demonstrating high within-sample accuracy.
@@
Table 2: Performance of DSP-based stutter removal methods using MFCC and LPC features across two stuttering types [10].
@@
For example, StutterNet, introduced by Sheikh et al., is a multi-class Time Delay Neural Network (TDNN). Because TDNNs use windows of time-delayed inputs, they are especially well-suited to temporal dependencies such as MFCC speech features. While StutterNet was only able to detect stutters, it is plausible that similar systems could be used to flag sections of stuttered audio for removal [11]. Table 3 summarizes representative DSP-based systems, their methodological choices, and reported performance.
@@
Table 3: Summary of studies on stutter detection, including datasets, feature extraction methods, detection approaches, stutter types analyzed, and best reported accuracies [10].
@@ 1.0.2 ASR & TTS-based Approaches
Mujtaba & Mahapatra describe fine-tuning Whisper-small, a state-of-the-art ASR model pretrained on fluent speech corpora. To allow for efficient training, low-rank adaptation (LoRA) was used as a form of parameter-efficient fine-tuning. Whisper-small was fine-tuned on ground truth transcripts provided by the FluencyBank and private HeardAI datasets. They reported that Whisper-Small achieved a WER of 33.88% without any training or fine-tuning. After fine-tuning Whisper-Small using the aforementioned LoRA methods, it achieved a WER of 9.39% [12].
Due to the adaptability of fine-tuned ASR models, there has been research into fine-tuning pre-trained ASR models for other speech impairments such as dysarthria. Wang et al. fine-tuned wav2vec2.0 and HuBERT ASR models on dysarthria datasets. They achieved the best WER of 16.53% by fine-tuning a pre-trained wav2vec2.0 ASR model on a dysarthria corpus augmented by a generative adversarial network. Even without any data augmentation, they obtained a WER of 22.25% on a fine-tuned wav2vec2.0 model and a WER of 21.10% on a fine-tuned HuBERT model [13].
@@ 1.0.3 Deep Learning & End-to-end Approaches
Deep learning approaches for speech recognition and speech conversion have been the subject of extensive investigation. Speech conversion focuses on modifying or transforming a speech signal, such as changing the speaker’s voice or style, while preserving the original linguistic content. In contrast, speech recognition involves accurately transcribing spoken language into text [14] [15] [16]. Deep learning approaches for speech processing typically involve training neural networks or other computational models, eliminating the need for manual feature engineering. Recent advancements in deep learning have popularized end-to-end models, systems that learn to perform the entire correction process directly from input data without requiring handcrafted features or intermediate steps.
@@
The popularity of end-to-end models stems from their ability to consolidate the entire training and inference pipeline into a single model optimized with one objective function that directly reflects the training goal. In contrast, a traditional DSP-based pipeline might first extract MFCC features from audio and then train a neural network to classify whether each frame contains a stutter. The neural network in this case is trained for frame-level accuracy in detecting stutters, which does not align with removing stutters at the speech sequence level [17] [18]. The lack of manual feature engineering in end-to-end models also allows for greater generalizability and adaptability to similar problems [19]. Finally, end-to-end models such as encoder-decoder and RNN-Transducer models are suitable for sequence-to-sequence tasks such as speech conversion, as they do not require alignment of acoustic sequences to ground truth transcripts and are very flexible with input and output sequence lengths [16].
For example, Toshniwal et al. developed an attention-based encoder-decoder model inspired by recurrent neural networks [20]. A speech encoder is built on a deep bidirectional long short-term memory (BiLSTM) network, which produces a sequence of abstract hidden representations. These hidden representations are passed into the character decoder, which is a single-layer LSTM that predicts the most likely letter uttered [20].
Wang [21] introduced an end-to-end encoder-decoder for dysarthric speech correction. Three multitask encoders were used: a content encoder to learn underlying semantic meaning, a prosody encoder to learn and correct audio features salient to dysarthria, and a speaker encoder to capture prosody and recreate the tone and voice of the original speaker. A decoder aggregates hidden representations from all three layers and generates acoustic features from which audio can be reconstructed using a vocoder [21].
@@ 2.1 Data
This research uses two datasets for training: Stuttering Events in Podcasts (SEP-28K) and LibriStutter. The SEP-28K dataset contains 23 hours of naturally occurring stuttering events, divided into 28,000 clips, each lasting 3 seconds [8]. While this dataset does not contain any ground truth for what the fluent speech should be, it includes labels classifying every stutter into the categories shown in Table 1. All audio recordings were taken from public podcasts with people who stutter at a standard sampling rate of 16 kHz. After all cleaning and processing, about 14 hours of raw audio data were left.
About 20 hours of artificially produced stuttered audio are included in the LibriStutter dataset. This corpus was created by splicing, cutting, duplicating, and performing other manipulations on the fluent LibriSpeech corpus to mimic the prosodic characteristics of stuttering [2]. SEP-28K does not include fluent reference transcripts, whereas LibriStutter contains paired fluent transcriptions that facilitate supervised training.
@@ 2.1.1 Data Cleaning and Preprocessing
The University College London Archive of Stuttered Speech (UCLASS) and the FluencyBank dataset are the two other well-known stuttering datasets [22] [23]. However, both datasets have some limitations. The UCLASS dataset does not provide fluent ”ground truth” transcripts for the stuttered speech, so there are no reference transcripts to fine-tune Whisper-Small on.
@@ 2.3 Whisper-Small Architecture
This study fine-tuned Whisper-Small, a lightweight derivative of the Whisper family of ASR models. Audio data from both datasets are in the amplitude domain, though Whisper-Small accepts a spectrogram in the frequency domain as input [24].
Whisper-Small’s WhisperFeatureExtractor uses a Short-Time Fourier Transform (STFT), computing the Fast Fourier Transform (FFT) on overlapping segments, or “windows,” of the audio as it slides across the entire signal. Then, the spectrogram is converted into a log-Mel spectrogram by mapping the frequencies from the “vanilla” spectrogram onto the Mel scale. This is done because human hearing does not perceive pitch in a linear manner; humans are more attuned to changes in lower pitches than in higher pitches. The linear frequency spectrogram is passed through a Mel filter bank, which applies a series of triangular filters to aggregate spectral energy within perceptually relevant frequency bands. Once a spectrogram is transformed into a log-Mel spectrogram, equal intervals on the Mel scale reflect equal perceived differences in pitch. The practical effect is that Mel spectrograms highlight the frequencies of human speech while minimizing the intensity of background noise, allowing the model to concentrate solely on the speech signals [25]. Figure 2 displays visualized spectrograms of the conversion process used to obtain a log-Mel spectrogram.
@@
Whisper-Small is a Transformer encoder-decoder model featuring 12 layers in both the encoder and decoder, a hidden width of 768 units, and 12 attention heads, amounting to around 244 million parameters in total. The encoder layers consist of a self-attention mechanism and fully connected hidden layers. The encoder generates a context vector, which is passed via cross-attention to each of the 12 decoder attention blocks. Each decoder layer contains self-attention, cross-attention, and a fully connected network [24]. At every time step, the autoregressive decoder is fed the aggregated output tokens from all previous time steps so that it can predict the most likely following token.
@@ 2.4 Whisper-Small Fine-tuning
Fine-tuning begins with Whisper-Small’s pretrained weights to take advantage of the accurate general transcription properties that Whisper is already trained for. The evaluation metric is Word Error Rate (WER), which is formally defined in Equation (1) [26].
@@ 2.5 StutterZero Model Architecture
This study introduces an autoregressive, end-to-end, multitask model named StutterZero. Inputs consist of 80-channel log-Mel spectrograms computed with a 50 ms window length and a 12.5 ms frame shift. The encoder converts the input log-Mel spectrogram into a higher-level representation called a context vector. The context vector is a numerical representation of features that the trained encoder determines to be relevant. While typical encoder-decoder models use one encoder and one decoder, this study proposes a multitask decoder in which two decoders are forced to predict different data types. During training, a multitask model minimizes a joint loss function, forcing the decoders to learn more generalized patterns that benefit all constituent losses [27]. The spectrogram decoder predicts the fluent spectrogram signal, while the transcript decoder predicts the grapheme being uttered.
@@ 2.5.1 StutterZero Encoder
The encoder consists of two convolutional blocks, each including a two-dimensional convolution layer with a 3×3 kernel, a 2×2 stride, and batch normalization. To effectively capture both spatial and temporal features of the input sequence, a convolution-augmented bidirectional long short-term memory network (ConvBiLSTM) is employed. Log-Mel spectrograms provide information about both the frequency content of a signal and the timing of its occurrence. Unlike traditional LSTMs, the ConvBiLSTM replaces standard matrix multiplication with a convolution operation, which promotes the learning of local spatial patterns in two-dimensional data. The network includes mechanisms analogous to the standard LSTM gates—input, forget, and output gates—as well as candidate cell states and hidden states. Each gate and state use convolutional kernels and trainable biases to process the current input in combination with the hidden state from the previous time step, enabling the model to capture complex spatiotemporal dependencies [28].

### v2

@@ 1 Introduction
Stuttering is a prevalent speech disorder that affects more than 70 million individuals worldwide [ghai2021]. It manifests as interruptions in speech flow, including unintentional repetitions, prolongations, and pauses. Table 1 presents the five main phonological classifications of stuttering and examples of their symptoms [kourkounakis2020a] [basak2023]. Such interruptions can significantly hinder effective communication, frequently causing social anxiety, depression, and a diminished quality of life [iverach2009]. In children, stuttering may lead to bullying and social isolation, exacerbating the emotional and psychological difficulties they encounter [iverach2016].
@@
Fluent speech plays a critical role in daily communication, both in interpersonal situations and when engaging with intelligent voice-based technologies. Through a quality-of-life survey, people who stutter (PWS) showed a statistically significant decrease in emotional health and social function metrics [kasbi2015]. Stuttering can also cause feelings of shame, fear, anxiety, and guilt [bloodstein2021]. These challenges are exacerbated by the growing dependence on intelligent voice assistants, such as Google Home, Amazon Echo, and Apple Siri, which are designed to process fluent speech. As a result, people who stutter frequently encounter recognition errors, limited functionality, and exclusion when interacting with voice-controlled technologies.
For more severe forms of stuttering, automatic speech recognition (ASR) models, which transcribe speech, may insert unintended words or even truncate the speech due to a blocking stutter. Lea et al. [lea2023] evaluated the speech of individuals with PWS using the Apple Speech framework, a production-level automatic speech recognition (ASR) system designed for fluent speakers. They reported a Word Error Rate (WER) of 19.8%, indicating that 19.8% of words in the ASR-generated transcript did not match the reference ground truth. They also reported a truncation rate of 23.8%, where 23.8% of utterances from these PWS were prematurely cut off [lea2023]. Addressing this challenge, this study develops deep learning models that convert stuttered speech into fluent audio, going beyond transcription to reconstruct corrected acoustic signals. By generating fluent speech that preserves semantic content and improves intelligibility, these models aim to enhance communication for PWS and increase accessibility in human-machine interactions.
@@ 1.0.1 Conventional Digital Signal Processing Approaches
DSP approaches utilize audio feature extraction methods to obtain condensed numerical features from complex audio signals. Then, a ruleset or a set of predetermined filters is applied to the features to determine which timeframes contain a stutter. Finally, these timeframes are cut out of the audio, removing the stutter. Some frequently utilized feature sets include Mel-Frequency Cepstral Coefficients (MFCCs), Linear Predictive Coding (LPC), and Linear Predictive Cepstral Coefficients (LPCCs) [sheikh2022]. For instance, MFCC features are generated by first applying the Fast Fourier Transform (FFT) to convert an audio sample from the amplitude domain to the frequency domain, generating the power spectrum. After that, the Mel filter bank is used to map the power spectrum to Mel frequencies, employing a set of nonlinear triangular filters. This aligns the intensity of frequencies to match the nonlinearity of human auditory perception. Finally, a Discrete Cosine Transform (DCT) is used to generate the cepstral coefficients through the decorrelation of features. This pipeline is applied to every window of audio, usually 20–50 ms long [sheikh2022] [kn2020].
K. N. et al. introduced a rule-based approach for stutter detection and removal by computing a “correlation factor” between adjacent audio windows using either MFCC or LPC features [kn2020]. A high correlation value, empirically determined to be 0.92, was used to identify repeated or prolonged segments, prompting the deletion of redundant frames. The unusually high correlation factor was used to detect repeated audio patterns, such as recurring phonemes or extended silences. However, this approach was evaluated exclusively on repetition and prolongation stutters, without accounting for other common disfluencies such as blocks or interjections. Additionally, the experimental scope of the study was confined to only 60 repetition and 70 prolongation events obtained from a limited selection of audio files. As a result, its generalizability across diverse speakers, accents, or spontaneous conversational settings remains uncertain [kn2020]. Table 2 summarizes the performance, demonstrating high within-sample accuracy.
@@
Table 2: Performance of DSP-based stutter removal methods using MFCC and LPC features across two stuttering types [kn2020].
@@
For example, StutterNet, introduced by Sheikh et al., is a multi-class Time Delay Neural Network (TDNN). Because TDNNs use windows of time-delayed inputs, they are especially well-suited to temporal dependencies such as MFCC speech features. While StutterNet was only able to detect stutters, it is plausible that similar systems could be used to flag sections of stuttered audio for removal [sheikh2021]. Table 3 summarizes representative DSP-based systems, their methodological choices, and reported performance.
@@
Table 3: Summary of studies on stutter detection, including datasets, feature extraction methods, detection approaches, stutter types analyzed, and best reported accuracies [kn2020].
@@ 1.0.2 ASR & TTS-based Approaches
Mujtaba & Mahapatra describe fine-tuning Whisper-small, a state-of-the-art ASR model pretrained on fluent speech corpora. To allow for efficient training, low-rank adaptation (LoRA) was used as a form of parameter-efficient fine-tuning. Whisper-small was fine-tuned on ground truth transcripts provided by the FluencyBank and private HeardAI datasets. They reported that Whisper-Small achieved a WER of 33.88% without any training or fine-tuning. After fine-tuning Whisper-Small using the aforementioned LoRA methods, it achieved a WER of 9.39% [mujtaba2025].
Due to the adaptability of fine-tuned ASR models, there has been research into fine-tuning pre-trained ASR models for other speech impairments such as dysarthria. Wang et al. fine-tuned wav2vec2.0 and HuBERT ASR models on dysarthria datasets. They achieved the best WER of 16.53% by fine-tuning a pre-trained wav2vec2.0 ASR model on a dysarthria corpus augmented by a generative adversarial network. Even without any data augmentation, they obtained a WER of 22.25% on a fine-tuned wav2vec2.0 model and a WER of 21.10% on a fine-tuned HuBERT model [wang2024a].
@@ 1.0.3 Deep Learning & End-to-end Approaches
Deep learning approaches for speech recognition and speech conversion have been the subject of extensive investigation. Speech conversion focuses on modifying or transforming a speech signal, such as changing the speaker’s voice or style, while preserving the original linguistic content. In contrast, speech recognition involves accurately transcribing spoken language into text [triantafyllopoulos2023] [walczyna2023] [wang2019]. Deep learning approaches for speech processing typically involve training neural networks or other computational models, eliminating the need for manual feature engineering. Recent advancements in deep learning have popularized end-to-end models, systems that learn to perform the entire correction process directly from input data without requiring handcrafted features or intermediate steps.
@@
The popularity of end-to-end models stems from their ability to consolidate the entire training and inference pipeline into a single model optimized with one objective function that directly reflects the training goal. In contrast, a traditional DSP-based pipeline might first extract MFCC features from audio and then train a neural network to classify whether each frame contains a stutter. The neural network in this case is trained for frame-level accuracy in detecting stutters, which does not align with removing stutters at the speech sequence level [chan2016] [graves2014]. The lack of manual feature engineering in end-to-end models also allows for greater generalizability and adaptability to similar problems [hannun2014]. Finally, end-to-end models such as encoder-decoder and RNN-Transducer models are suitable for sequence-to-sequence tasks such as speech conversion, as they do not require alignment of acoustic sequences to ground truth transcripts and are very flexible with input and output sequence lengths [wang2019].
For example, Toshniwal et al. developed an attention-based encoder-decoder model inspired by recurrent neural networks [toshniwal2017]. A speech encoder is built on a deep bidirectional long short-term memory (BiLSTM) network, which produces a sequence of abstract hidden representations. These hidden representations are passed into the character decoder, which is a single-layer LSTM that predicts the most likely letter uttered [toshniwal2017].
Wang [wang2022] introduced an end-to-end encoder-decoder for dysarthric speech correction. Three multitask encoders were used: a content encoder to learn underlying semantic meaning, a prosody encoder to learn and correct audio features salient to dysarthria, and a speaker encoder to capture prosody and recreate the tone and voice of the original speaker. A decoder aggregates hidden representations from all three layers and generates acoustic features from which audio can be reconstructed using a vocoder [wang2022].
@@ 2.1 Data
This research uses two datasets for training: Stuttering Events in Podcasts (SEP-28K) and LibriStutter. The SEP-28K dataset contains 23 hours of naturally occurring stuttering events, divided into 28,000 clips, each lasting 3 seconds [lea2023]. While this dataset does not contain any ground truth for what the fluent speech should be, it includes labels classifying every stutter into the categories shown in Table 1. All audio recordings were taken from public podcasts with people who stutter at a standard sampling rate of 16 kHz. After all cleaning and processing, about 14 hours of raw audio data were left.
About 20 hours of artificially produced stuttered audio are included in the LibriStutter dataset. This corpus was created by splicing, cutting, duplicating, and performing other manipulations on the fluent LibriSpeech corpus to mimic the prosodic characteristics of stuttering [kourkounakis2020a]. SEP-28K does not include fluent reference transcripts, whereas LibriStutter contains paired fluent transcriptions that facilitate supervised training.
@@ 2.1.1 Data Cleaning and Preprocessing
The University College London Archive of Stuttered Speech (UCLASS) and the FluencyBank dataset are the two other well-known stuttering datasets [zotero-item-1105] [bernsteinratner2018]. However, both datasets have some limitations. The UCLASS dataset does not provide fluent ”ground truth” transcripts for the stuttered speech, so there are no reference transcripts to fine-tune Whisper-Small on.
@@ 2.3 Whisper-Small Architecture
This study fine-tuned Whisper-Small, a lightweight derivative of the Whisper family of ASR models. Audio data from both datasets are in the amplitude domain, though Whisper-Small accepts a spectrogram in the frequency domain as input [radford2022].
Whisper-Small’s WhisperFeatureExtractor uses a Short-Time Fourier Transform (STFT), computing the Fast Fourier Transform (FFT) on overlapping segments, or “windows,” of the audio as it slides across the entire signal. Then, the spectrogram is converted into a log-Mel spectrogram by mapping the frequencies from the “vanilla” spectrogram onto the Mel scale. This is done because human hearing does not perceive pitch in a linear manner; humans are more attuned to changes in lower pitches than in higher pitches. The linear frequency spectrogram is passed through a Mel filter bank, which applies a series of triangular filters to aggregate spectral energy within perceptually relevant frequency bands. Once a spectrogram is transformed into a log-Mel spectrogram, equal intervals on the Mel scale reflect equal perceived differences in pitch. The practical effect is that Mel spectrograms highlight the frequencies of human speech while minimizing the intensity of background noise, allowing the model to concentrate solely on the speech signals [stevens1937]. Figure 2 displays visualized spectrograms of the conversion process used to obtain a log-Mel spectrogram.
@@
Whisper-Small is a Transformer encoder-decoder model featuring 12 layers in both the encoder and decoder, a hidden width of 768 units, and 12 attention heads, amounting to around 244 million parameters in total. The encoder layers consist of a self-attention mechanism and fully connected hidden layers. The encoder generates a context vector, which is passed via cross-attention to each of the 12 decoder attention blocks. Each decoder layer contains self-attention, cross-attention, and a fully connected network [radford2022]. At every time step, the autoregressive decoder is fed the aggregated output tokens from all previous time steps so that it can predict the most likely following token.
@@ 2.4 Whisper-Small Fine-tuning
Fine-tuning begins with Whisper-Small’s pretrained weights to take advantage of the accurate general transcription properties that Whisper is already trained for. The evaluation metric is Word Error Rate (WER), which is formally defined in Equation (1) [hunt1990].
@@ 2.5 StutterZero Model Architecture
This study introduces an autoregressive, end-to-end, multitask model named StutterZero. Inputs consist of 80-channel log-Mel spectrograms computed with a 50 ms window length and a 12.5 ms frame shift. The encoder converts the input log-Mel spectrogram into a higher-level representation called a context vector. The context vector is a numerical representation of features that the trained encoder determines to be relevant. While typical encoder-decoder models use one encoder and one decoder, this study proposes a multitask decoder in which two decoders are forced to predict different data types. During training, a multitask model minimizes a joint loss function, forcing the decoders to learn more generalized patterns that benefit all constituent losses [zhang2021]. The spectrogram decoder predicts the fluent spectrogram signal, while the transcript decoder predicts the grapheme being uttered.
@@ 2.5.1 StutterZero Encoder
The encoder consists of two convolutional blocks, each including a two-dimensional convolution layer with a 3×3 kernel, a 2×2 stride, and batch normalization. To effectively capture both spatial and temporal features of the input sequence, a convolution-augmented bidirectional long short-term memory network (ConvBiLSTM) is employed. Log-Mel spectrograms provide information about both the frequency content of a signal and the timing of its occurrence. Unlike traditional LSTMs, the ConvBiLSTM replaces standard matrix multiplication with a convolution operation, which promotes the learning of local spatial patterns in two-dimensional data. The network includes mechanisms analogous to the standard LSTM gates—input, forget, and output gates—as well as candidate cell states and hidden states. Each gate and state use convolutional kernels and trainable biases to process the current input in combination with the hidden state from the previous time step, enabling the model to capture complex spatiotemporal dependencies [shi2015].